UT 20131113 LUCI2 Nighttime Commissioning

Observer: WSeifert, JHeidt, PBuschkamp, APramskiy, KPolsterer, MEdwards, DThompson
Telescope Operator: SAllanson
LBTO Support: TSargent

Plan for the night

Lots of clouds around at the beginning of the night, we start with closed-dome calibs

Summary

Clouded out, no obs done

Could derive flexure compensation data in 6 combos, have now a full set of flexure data

Preparations

Details

Files are luci2.20131112.iii

Afternoon DThompson took a set of 20 times 2.7sec darks for RON, files #11-30

Analysis by Dave: RON determined by spatial method (using difference of two frames) => 12.5 e-, temporal 12.3 e-

He had problems with GEIRS (eg came up with IDLE-wait), took some time to get rid of it

0:39 Sunset, Weather: mostly cloudy, Temp = +3deg, Hum 75%, strong winds (13m/s)

Flexure data

Script flexure_minus1_with_delay_N375/180, t_exp=5sec for N1.8, 20sec for N375-camera
Ima-mode: clear+K-filter, spec-mode: Ar/clear+clear, Lamda_cen=2.25,1.93,2.2mue (G150,G200/G210)

01:00 N1.8, imaging, files #52-54, stopped after 3rd exp., GEIRS error

01:07 N1.8, imaging, #55-70, stopped after 15th exp., restarted GEIRS

01:26 N1.8, imaging, files #72-73, again problems with GEIRS,

02:05 Sky is a bit more clear now, but both - hum and wind close to the limits => stay closed
Problems with GEIRS solved (driver)

02:06 Ele=90, N1.8, ima, files #74-110
02:31 Ele=55, N1.8, ima, files #111-147
02:57 Ele=20, N1.8, ima, files #148-184, Dark 185

03:25 Ele=90, N1.8, G200, got stuck at ROT=270, restart, got GEIRS-error, restart, files #199-235
04:39 Ele=55, N1.8, G200, files #236-272
05:10 Ele=20, N1.8, G200, files #273-309, Dark 310

06:30 Sky partly clear now, but wind and hum above limits

06:39 Ele=90, N1.8, G150, execution panel hangs after 9th exp., restart GEIRS & SW, files #320-356
06:58 Ele=55, N1.8, G150, files #357-393
07:22 Ele=20, N1.8, G150, files #394-419, hangs after 20th exposure), restart GEIRS & SW, files #420-456
There seems to be something fundamentally wrong, the ROM-GUI and/or GEIRS hang up frequently this night

08:44 Ele=90, N3.75, G210, files #459-495
09:15 Ele=55, N3.75, G210, files #496-532
09:48 Ele=20, N3.75, G210, files #533-569, Dark 570

10:00 Weather: sitting in fog and strong winds

10:20 Ele=90, N3.75, G200, files #571-607
10:55 Ele=55, N3.75, G200, files #608-644
11:27 Ele=20, N3.75, G200, files #645-681

12:02 Ele=90, N3.75, G150, files #683-719
12:34 Ele=55, N3.75, G150, files #720-756
13:08 Ele=20, N3.75, G150, files #757-761, got again stuck at 761, exposing forever,
fixed by Alexander (GEIRS socket problem), #763-799

14:09 Finito

-- JochenHeidt - 13 Nov 2013
Topic revision: r7 - 05 Dec 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback